### v2

# Decoding the string in terms of holographic quantum maps

Abstract
It has recently been shown that the Nambu-Goto equation for a string emerges from the junction conditions in three-dimensional gravity. Holographically, gravitational junctions are dual to interfaces in conformal field theory. We demonstrate that each stringy mode of the junction corresponds to a universal $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ quantum map between in and out Hilbert spaces of excitations scattered at the interface, and also a universal $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ quantum map relating the excitations on both sides. These quantum maps generalize those realized by defect operators, preserve the conformal boundary condition at the interface, and realize a tunable energy transmitter.
Introduction:- The holographic duality states that quantum gravity, particularly a superstring theory, can be reformulated in terms of a nongravitational quantum field theory living at the boundary of spacetime [1, 2, 3]. There has been a lot of progress in our understanding of how the bulk spacetime and semi-classical bulk effective theory consisting of Einstein’s gravity coupled to a few fields can be reconstructed from the boundary quantum field theory in the large N and strong coupling limit [4, 5, 6, 7]. The reconstruction of dynamical extended objects (branes) of the gravitational theory that are necessary for its non-perturbative completion, is therefore a problem of fundamental importance.
Gravitational junctions between three-dimensional asymptotically anti-de-Sitter (AdS3) spacetimes model properties of conformal interfaces in dual two-dimensional conformal field theories (CFTs) [8, 9, 10, 11], which appear also in many condensed matter systems, e.g. in quantum wire junctions [12, 13, 14], and defects in spin systems [15] and quantum Hall systems [16, 17, 18].
It has been shown that the Nambu-Goto equation for the string emerges from the gravitational junction conditions in three dimensions [19]. In this letter, we report how these stringy modes of gravitational junctions correspond to universal $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ quantum maps from the in to out Hilbert spaces of excitations scattered at the dual interface. Equivalently, the correspondence can be formulated via universal $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ quantum maps relating the Hilbert spaces of the two CFTs that straddle the interface. These maps generalize the $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ and $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ maps realized by a defect operator while preserving the conformal boundary condition at the interface. We also prove the universality of the quantum maps corresponding to the holographic defect operator.
The conformal interface:- The non-vanishing components of the energy-momentum tensor $T_{\mu\nu}$ in any state of a two-dimensional CFT are $T_{\pm\pm}=T_{\mu\nu}n_{\pm}^{\mu}n_{\pm}^{\nu}$, with $n_{\pm}^{\mu}$ being the two future directed null vectors. The integration of $T_{++}$ and $T_{--}$ at past null infinity define the operators $\mathcal{E}_{\pm}$ which measure the left-moving and right-moving incoming energy fluxes, respectively. A conformal interface is a gluing of two CFTs, namely CFTL and CFTR on the left and right, respectively, such that scale invariance is preserved. The interface at $x=0$ is represented by an operator insertion $I_{L,R}$ which satisfies
$(L_{n,+}^{L}-L_{-n,-}^{L})I_{L,R}=I_{L,R}(L_{n,+}^{R}-L_{-n,-}^{R}),$ (1)
[…]
Either way, the interface represents a linear map $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ from states in CFTL to states in CFTR.
The interface can also be viewed as a $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ map, e.g. the $S$ matrix relating the incoming right-moving and left-moving energy excitations on the left and right sides of the interface, respectively, and the outgoing left-moving and right-moving energy excitations on the respective left and right sides (see Fig. 1). When both CFTs have the same central charge $c$, this $S$ matrix is
$\begin{pmatrix}\langle\mathcal{E}_{+}^{L}\rangle\\ \langle\mathcal{E}_{-}^{R}\rangle\end{pmatrix}=S\begin{pmatrix}\langle\mathcal{E}_{-}^{L}\rangle\\ \langle\mathcal{E}_{+}^{R}\rangle\end{pmatrix},\quad S=\begin{pmatrix}\frac{2}{2+\lambda}&\frac{\lambda}{2+\lambda}\\ \frac{\lambda}{2+\lambda}&\frac{2}{2+\lambda}\end{pmatrix}$, (3)
with $\lambda=\frac{2(c-c_{LR})}{c_{LR}}$, where $c_{LR}$ is the coefficient that appears in the two-point function $\langle T_{++}^{L}(x)T_{++}^{R}(y)\rangle_{I}$ in presence of the defect operator [21]. Clearly, $S$ is simply the independent transmission/reflection of the incoming energies from the left and right with equal transmission coefficients
$\mathbb{T}=\frac{2}{2+\lambda},$ (4)
and equal reflection coefficients $\mathbb{R}=1-\mathbb{T}$, implying energy conservation ($\langle\mathcal{E}_{-}^{L}\rangle+\langle\mathcal{E}_{+}^{R}\rangle=\langle\mathcal{E}_{-}^{R}\rangle+\langle\mathcal{E}_{+}^{L}\rangle$), which follows from the conformal boundary condition (1). Remarkably, $S$ is universal, i.e. independent of how the energy fluxes are created [21]. Crucially, an overall (diagonal) conformal transformation on CFTL and CFTR does not affect $S$. The achronal ANEC (averaged null energy condition) implies that $c_{LR}\leq c$ [21], i.e. $\lambda\geq 0$.
It is possible to generalize the conformal interface with a half-sided conformal transformation in which the same conformal transformation is applied to the left and right movers on either the left or the right side. Such a transformation preserves the interface at $x=0$ and the $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ map is then simply composed with a conformal transformation on CFTR, which results in a redefinition of the $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ map. We will discuss these explicitly later.
Gravitational two-way junction:- The conformal interface is holographically dual to a gravitational two-way junction. Consider two three-dimensional locally AdS3 manifolds $\mathcal{M}_{1,2}$. Each of these is split into two parts $\mathcal{M}_{i\alpha_{i}}$, $i=1,2$, $\alpha_{i}=L,R$, by co-dimension-1 hypersurfaces $\Sigma_{1,2}$. A gravitational junction $\Sigma$ involves the joining of one fragment each of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. Here we will glue $\mathcal{M}_{1L}$ and $\mathcal{M}_{2R}$. The full spacetime $\widetilde{\mathcal{M}}$ is formed by the gluing of $\mathcal{M}_{1L}$ and $\mathcal{M}_{2R}$ at $\Sigma$ by identifying points on $\Sigma_{1,2}$, which are the images of $\Sigma$ in $\mathcal{M}_{1,2}$. This identification of points and the embeddings of $\Sigma_{1,2}$ in $\mathcal{M}_{1,2}$ should satisfy the gravitational junction conditions [22].
Figure 1: A two-way gravitational junction formed by gluing two AdS3 manifolds by identifying points $P_{1,2}$ on the (gray) hypersurfaces. The incoming and outgoing plane wave amplitudes are indicated in the boundary CFT by arrows. The $a_{n}$ are classical string excitations.
Let $\mathcal{M}_{1,2}$ have the coordinates $t_{1,2},x_{1,2},z_{1,2}$, with $x_{1,2}$ the coordinates transverse to $\Sigma_{1,2}$ and $z_{1,2}$ the radial coordinates. The embeddings of $\Sigma_{1,2}$ are specified by two functions $f_{1,2}$
$\Sigma_{1,2}:\quad x_{1,2}=f_{1,2}(t_{1,2},z_{1,2}).$ (5)
The freedom of choice of the coordinates of $\Sigma$ is fixed by defining the worldsheet coordinates $\tau,\sigma$ at a point $P$ on $\Sigma$ as
$\tau(P)=\frac{t_{1}(P_{1})+t_{2}(P_{2})}{2},\quad\sigma(P)=\frac{z_{1}(P_{1})+z_{2}(P_{2})}{2},$ (6)
where $P_{1,2}$ are the points on $\Sigma_{1,2}$ that are identified with $P$ (see Fig. 1). The following four variables, which are functions of $\tau,\sigma$, completely specify the junction
$\tau_{d}=\frac{t_{2}-t_{1}}{2},\quad\sigma_{d}=\frac{z_{2}-z_{1}}{2},\quad x_{s,d}=\frac{f_{2}\pm f_{1}}{2}.$ (7)
The junction conditions can be obtained from the action
$S_{grav}=\frac{1}{16\pi G_{N}}\int_{\mathcal{\widetilde{\mathcal{M}}}}d^{3}x\sqrt{-g}(R-2\Lambda)+T_{0}\int_{\Sigma}{\rm d}\tau{\rm d}\sigma\,\sqrt{-\gamma}$
$\quad+{\rm GHY\ terms},$ (8)
where the metric is the only degree of freedom. Above GHY are the Gibbons-Hawking-York (GHY) boundary terms and $T_{0}$ is the tension. This action is defined assuming that the induced metrics $\gamma_{1,2}$ on $\Sigma_{1,2}$ are identical, which defines the worldsheet metric as
$\gamma_{\mu\nu}(\tau,\sigma)=\gamma_{1,\mu\nu}(\tau,\sigma)=\gamma_{2,\mu\nu}(\tau,\sigma).$ (9)
Varying (Decoding the string in terms of holographic quantum maps) away from the junction implies that the manifold is Einstein. At the junction, we have
$\left(K_{i,\mu\nu}-K_{i}\,\gamma_{i,\mu\nu}\right)|_{\rm disc}=8\pi G_{N}T_{0}\gamma_{\mu\nu},$ (10)
where $K_{i,\mu\nu}$ is the extrinsic curvature of $\Sigma_{i}$ in $\mathcal{M}_{i\alpha_{i}}$, $K_{i}=K_{i,\mu\nu}\gamma^{\mu\nu}$ and $|_{\rm disc}$ denotes the discontinuity. The bulk diffeomorphism symmetry implies that the left hand side of (10) is conserved. We therefore obtain only one independent equation from (10), which together with (9) yields four equations for the four unknown functions (7).
It has been shown in [19], when $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are copies of an Einstein manifold $\mathcal{M}$, that the general solutions of the junction conditions are in one-to-one correspondence with solutions of the non-linear Nambu-Goto equation for a worldsheet in $\mathcal{M}$. Specifically, the hypersurface
$\Sigma_{NG}:t=\tau,\quad z=\sigma,\quad x=x_{s}(\tau,\sigma),$
corresponds to a solution of the non-linear Nambu-Goto equation in $\mathcal{M}$ when the tension $T_{0}$ vanishes, while $\tau_{d}$, $\sigma_{d}$ and $x_{d}$ are fixed completely by $x_{s}$ up to six rigid parameters related to worldsheet and spacetime isometries, which are irrelevant for the present paper.
Here we will focus on locally AdS3 Bañados metrics [23]
$ds^{2}=\frac{dz^{2}}{z^{2}}+2dtdx\left(\mathcal{L}_{+}(x^{+})-\mathcal{L}_{-}(x^{-})\right)\\ -\frac{dt^{2}}{z^{2}}\left(1-z^{2}\mathcal{L}_{+}(x^{+})\right)\left(1-z^{2}\mathcal{L}_{-}(x^{-})\right)\\ +\frac{dx^{2}}{z^{2}}\left(1+z^{2}\mathcal{L}_{+}(x^{+})\right)\left(1+z^{2}\mathcal{L}_{-}(x^{-})\right),$ (11)
where $x^{\pm}=t\pm x$, and we have set $\Lambda=-1$. We assume that $\mathcal{M}_{1,2}$ have the above metrics with
$\mathcal{L}_{\pm}^{(1)}(x_{1}^{\pm})=L^{L}_{\pm}e^{i\omega x_{1}^{\pm}},\quad\mathcal{L}^{(2)}_{\pm}(x_{2}^{\pm})=L^{R}_{\pm}e^{i\omega x_{2}^{\pm}}.$ (12)
Furthermore, we assume that $L^{L,R}_{\pm}=\mathcal{O}(\epsilon)$ are small amplitudes. Our analysis will be linear and the above plane waves can be superposed to form wavepackets. At $\mathcal{O}(\epsilon^{0})$ we have the exact solution
$\tau_{d}=0,\quad\sigma_{d}=0,\quad x_{s}=0,\quad x_{d}=-\frac{\lambda\sigma}{\sqrt{4-\lambda^{2}}},$ (13)
where $\lambda=8\pi G_{N}T_{0}$ and $0\leq\lambda\leq 2$. The equations (9) and (10) can be solved perturbatively in $\epsilon$. At $\mathcal{O}(\epsilon)$ we get the following equation
$16\sigma\partial^{2}_{\tau}x_{s}+4(4-\lambda^{2})\left(2\partial_{\sigma}x_{s}-\sigma\partial_{\sigma}^{2}x_{s}\right)=-ie^{i\omega\tau}(-4+\lambda^{2})\sigma^{3}\omega\left((L_{-}^{R}-L_{+}^{L})e^{\frac{i\lambda\omega\sigma}{\sqrt{4-\lambda^{2}}}}+(L_{-}^{L}-L_{+}^{R})e^{-\frac{i\lambda\omega\sigma}{\sqrt{4-\lambda^{2}}}}\right),$ (14)
which is the just the linearized Nambu-Goto (NG) equation in empty AdS with sources proportional to $(L_{-}^{L}-L_{+}^{R})$ and $(L_{-}^{R}-L_{+}^{L})$ when $\lambda\to 0$. This implies that the correspondence between the Nambu-Goto equation and junction conditions shown in [19] generalizes (with sources) even when the backgrounds on both sides depart from each other. We obtain the following solution to (14)
$x_{s}$ $=\epsilon e^{i\omega\tau}\Bigg(\frac{-i(4-\lambda^{2})^{1/4}\left(\sin\left(\frac{2\sigma\omega}{\sqrt{4-\lambda^{2}}}\right)\left(A_{1}\sqrt{4-\lambda^{2}}+2A_{2}\sigma\omega\right)+\cos\left(\frac{2\sigma\omega}{\sqrt{4-\lambda^{2}}}\right)\left(A_{2}\sqrt{4-\lambda^{2}}-2A_{1}\sigma\omega\right)\right)}{2\sqrt{\pi}(\omega)^{3/2}}$
$\quad\quad\qquad+\frac{(L^{R}_{-}-L^{L}_{+})e^{\frac{i\lambda\sigma\omega}{\sqrt{4-\lambda^{2}}}}\left(2\sqrt{4-\lambda^{2}}\lambda\sigma\omega-i\left(\lambda^{2}-4\right)\left(\sigma^{2}\omega^{2}+2\right)\right)}{4\left(\lambda^{2}-4\right)\omega^{3}}$
$\quad\quad\qquad+\frac{(L^{L}_{-}-L^{R}_{+})e^{-\frac{i\lambda\sigma\omega}{\sqrt{4-\lambda^{2}}}}\left(-2\sqrt{4-\lambda^{2}}\lambda\sigma\omega-i\left(\lambda^{2}-4\right)\left(\sigma^{2}\omega^{2}+2\right)\right)}{4\left(\lambda^{2}-4\right)\omega^{3}}\Bigg),$ (15)
where the first line is the solution of the source-free (homogeneous) equation (14). Imposing ingoing boundary conditions at the Poincaré horizon [24, 25, 26] we obtain the coefficients $A_{1}=A_{nn}+A_{n}$ and $A_{2}=iA_{nn}$. Here $A_{nn}$ corresponds to a non-normalizable mode of the homogeneous NG equation, which is the causal response to bulk perturbations that travels from the boundary towards the Poincaré horizon. $A_{n}$ is an intrinsic normalizable (stringy) mode of the homogeneous NG equation. Both $A_{nn}$ and $A_{n}$ are determined by initial and boundary conditions as usual in Lorentzian holographic duality. Explicit solutions for the other variables are in the End-matter.
The energy-momentum tensors on both sides of the dual CFT interface can be extracted using holographic renormalization [27, 28]. Explicitly,
$\braket{{T}^{L}_{\pm}(x_{1}^{\pm})}=\frac{c\epsilon}{12\pi}e^{i\omega{x}_{1}^{\pm}}L_{\pm}^{L},$ (16)
$\braket{{T}^{R}_{\pm}(x_{2}^{\pm})}=\frac{c\epsilon}{12\pi}e^{i\omega{x}_{2}^{\pm}}L_{\pm}^{R}.$ (17)
The Dirichlet boundary conditions corresponding to the interface at $x_{1}=x_{2}=0$ impose $\lim_{\sigma\to 0}x_{d}=\lim_{\sigma\to 0}x_{s}=0$. From $\lim_{\sigma\to 0}x_{d}=0$ we obtain
$L_{-}^{L}+L_{+}^{R}=L_{-}^{R}+L_{+}^{L}.$ (18)
This is the conservation of energy at the interface, or equivalently the conformal boundary condition. Eq. (18) can be solved by
$L_{+}^{L}=\mathcal{T}_{L}L_{+}^{R}+(1-\mathcal{T}_{R})L_{-}^{L},\quad L_{-}^{R}=\mathcal{T}_{R}L_{-}^{L}+(1-\mathcal{T}_{L})L_{+}^{R},$ (19)
where $\mathcal{T}_{L,R}$ and $1-\mathcal{T}_{L,R}$ are the transmission and reflection coefficients for the left and right-movers, respectively. From $\lim_{\sigma\to 0}x_{s}=0$, we obtain that
$A_{nn}=i\frac{2\sqrt{\pi}(L_{-}^{L}\mathcal{T}_{R}-L_{+}^{R}\mathcal{T}_{L})}{\left(4-\lambda^{2}\right)^{3/4}\omega^{3/2}}.$ (20)
The string as a holographic quantum map:- In [29], it was shown that the transmission/reflection coefficients of the CFT interface can be derived by keeping the incoming energy flux only on one side, considering the causal response on the worldsheet with $A_{n}=0$ and using the Dirichlet boundary conditions on the other variables $\tau_{d}$ and $\sigma_{d}$. However, the gravitational problem is fundamentally non-linear, and therefore assuming independent and equal transmission from the left and right sides affect the higher-order perturbative expansions. Furthermore, in order to prove the universality of the scattering process in the pure gravity setup, it is necessary to show that it is independent of the modification of the background geometry (which is assumed to be the empty AdS3 space dual to the CFT vacuum) on both sides. While decoding the stringy normalizable mode $A_{n}$ in terms of quantum maps, we derive the matrix $S$ given by (3) for arbitrary input energy fluxes and show that it is independent of the background.
The key point is that, instead of setting Dirichlet boundary conditions for $\tau_{d}$ and $\sigma_{d}$, we should use relative conformal transformations to set up continuous coordinates across the dual interface at the boundary. The scattering process is realized in these coordinates in the dual theory. Let us define
$\lim_{\sigma\rightarrow 0}\tau_{d}(\tau,\sigma)=\epsilon\mathbb{t}_{d}(\tau)+\mathcal{O}(\epsilon^{2})$ (21)
so that at the boundary $t_{1,2}(\tau)=\tau\mp\epsilon\mathbb{t}_{d}(\tau)+\mathcal{O}(\epsilon^{2})$. In agreement with [19], we find that the relative time reparametrization $\mathbb{t}_{d}(\tau)$ encodes the normalizable stringy mode as both $\mathbb{t}_{d}(\tau)$ and $\lim_{\sigma\to 0}\frac{\sigma_{d}}{\sigma}$ are proportional to
$L_{-}^{L}(-2+\mathcal{T}_{R}(2+\lambda))+L_{+}^{R}(-2+\mathcal{T}_{L}(2+\lambda))+2a_{n},$ (22)
where $a_{n}=\frac{iA_{n}\lambda(4-\lambda^{2})^{3/4}\omega^{3/2}}{4\sqrt{\pi}}$.
The discontinuity in the time coordinates ($t_{1}$ and $t_{2}$) at the interface located at $x_{1}=x_{2}=0$ can be undone using separate conformal transformations on the two sides which involve coordinate transformations
$\widetilde{t}_{1,2}=\frac{1}{2}(\mathbb{h}_{1,2}^{-1}(t_{1,2}+x_{1,2})+\mathbb{h}^{-1}_{1,2}(t_{1,2}-x_{1,2})),$
$\widetilde{x}_{1,2}=\frac{1}{2}(\mathbb{h}_{1,2}^{-1}(t_{1,2}+x_{1,2})-\mathbb{h}_{1,2}^{-1}(t_{1,2}-x_{1,2})),$ (23)
and the associated Weyl transformations that brings the metric on both sides back to the Minkowski form. The most general choices of $\mathbb{h}_{1,2}$ at the linear order are
$\mathbb{h}_{2}(\tau)=\tau+\epsilon\alpha\mathbb{t}_{d}(\tau),\quad\mathbb{h}_{1}(\tau)=\tau+\epsilon(\alpha-2)\mathbb{t}_{d}(\tau),$ (24)
where $\alpha$ is an arbitrary constant parameter which acts as an overall (diagonal) conformal transformation that changes the background state (see below). It is easy to see using (24) that (Decoding the string in terms of holographic quantum maps) preserves the interface at $\tilde{x}_{1,2}=0$ where $\tilde{t}_{2}=\tilde{t}_{1}$. We therefore obtain continuous coordinates and metric across the interface. These conformal transformations can be uplifted to bulk diffeomorphisms ($X^{\mu}(\tau,\sigma)\rightarrow\widetilde{X}^{\mu}(\tau,\sigma)$ with $X^{\mu}$ denoting bulk coordinates on both sides [28, 26, 30], and both the induced metric and the extrinsic curvatures of $\Sigma_{1,2}$ remain invariant under these bulk diffeomorphisms, producing an equivalent solution of the junction conditions.
Under the conformal transformations, the energy-momentum tensors on both sides of the interface transform as
[…]
$=\frac{c\epsilon}{12\pi}e^{i\omega\tilde{x}_{2}^{\pm}}\tilde{L}_{\pm}^{L,R}+\mathcal{O}(\epsilon^{2}),$ (25)
where
$\tilde{L}_{+}^{L}$ $=\frac{L_{+}^{R}}{4}\left(4+2(-1+\mathcal{T}_{L})\alpha+\mathcal{T}_{L}(-2+\alpha)\lambda\right)$
$\quad+\frac{L_{-}^{L}}{4}\left(2\alpha+\mathcal{T}_{R}(2\lambda-\alpha(2+\lambda))\right)+\frac{a_{n}(\alpha-2)}{2},$ (26)
$\tilde{L}_{-}^{L}$ $=\frac{L_{+}^{R}}{4}(\alpha-2)(-2+\mathcal{T}_{L}(2+\lambda))$
$\quad+\frac{L_{-}^{L}}{4}\left(2\alpha-\mathcal{T}_{R}(-2+\alpha)(2+\lambda)\right)+\frac{a_{n}(\alpha-2)}{2},$ (27)
$\tilde{L}_{+}^{R}$ $=\frac{L_{+}^{R}}{4}\left(4+\alpha(-2+\mathcal{T}_{L}(2+\lambda))\right)$
$\quad-\frac{L_{-}^{L}}{4}\alpha\left(-2+\mathcal{T}_{R}(2+\lambda)\right)+\frac{\alpha a_{n}}{2},$ (28)
$\tilde{L}_{-}^{R}$ $=\frac{L_{+}^{R}}{4}\left(2(-1+\mathcal{T}_{L})(-2+\alpha)+\mathcal{T}_{L}\alpha\lambda\right)$
$+\frac{L_{-}^{L}}{4}\left(2\alpha+\mathcal{T}_{R}(4-\alpha(2+\lambda))\right)+\frac{\alpha a_{n}}{2}.$ (29)
These transformations are reproduced in the holographic renormalization procedure of extracting the dual energy-momentum tensors via the bulk diffeomorphisms which uplift the corresponding conformal transformations [30]. The transformed amplitudes also satisfy energy conservation (and thus also conformal boundary condition) at linear order as
$\tilde{L}_{-}^{L}+\tilde{L}_{+}^{R}=\tilde{L}_{-}^{R}+\tilde{L}_{+}^{L}.$ (30)
As evident from (26), (27), (28) and (29), the four physical energy fluxes $\tilde{L}_{\pm}^{L,R}$ depend on the six parameters $\alpha$, $a_{n}$, $_{+}^{R}, LL$$_{-}^{L}$, $\mathcal{T}_{L}$ and $\mathcal{T}_{R}$. Remarkably, for arbitrary values of the six parameters, we obtain that the physical energy fluxes satisfy
$\tilde{L}_{+}^{L}$ $=\frac{2}{2+\lambda}\left(\tilde{L}_{+}^{R}-\frac{2a_{n}}{2-\lambda}\right)+\frac{\lambda}{2+\lambda}\left(\tilde{L}_{-}^{L}+\frac{2a_{n}}{2-\lambda}\right),$
$\tilde{L}_{-}^{R}$ $=\frac{\lambda}{2+\lambda}\left(\tilde{L}_{+}^{R}-\frac{2a_{n}}{2-\lambda}\right)+\frac{2}{2+\lambda}\left(\tilde{L}_{-}^{L}+\frac{2a_{n}}{2-\lambda}\right),$ (31)
This implies that the dual interface acts as a $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ quantum map given by $\mathcal{S}\circ\mathcal{D}$ where $S$ is the matrix (3) and $\mathcal{D}$ is a redistribution of energy among the two incoming energy fluxes, which can be realized by the linearized conformal transformation given by
$g_{\pm}(x^{\pm})=x^{\pm}\mp\epsilon\frac{e^{i\omega x^{\pm}}}{\omega^{3}}\frac{4ia_{n}}{2-\lambda}+\mathcal{O}(\epsilon^{2}).$ (32)
The physical fluxes (26), (27), (28) and (29) also satisfy
$\tilde{L}_{+}^{L}$ $=\frac{2}{2+\lambda}\tilde{L}_{+}^{R}+\frac{\lambda}{2+\lambda}\tilde{L}_{-}^{L}-\frac{2a_{n}}{2+\lambda},$
$\tilde{L}_{-}^{R}$ $=\frac{\lambda}{2+\lambda}\tilde{L}_{+}^{R}+\frac{2}{2+\lambda}\tilde{L}_{-}^{L}+\frac{2a_{n}}{2+\lambda},$ (33)
so that the $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ map can be rewritten in the form $\widetilde{\mathcal{D}}\circ\mathcal{S}$ where $\widetilde{\mathcal{D}}$ is a redistribution of energy among the two outgoing energy fluxes which can also be realized by a linearized conformal transformation. The $\mathcal{H}_{in}\rightarrow\mathcal{H}_{out}$ map in both forms, namely $\mathcal{S}\circ\mathcal{D}$ and $\widetilde{\mathcal{D}}\circ\mathcal{S}$, is independent of the background state (represented by the choice of $\alpha$) and reduces to the universal scattering matrix $S$ given by (3) in absence of the stringy mode $a_{n}$, which otherwise gives rise to universal energy redistribution among the incoming or outgoing energy fluxes.
We also note from (26), (27), (28) and (29) that
$\tilde{L}_{+}^{R}$ $=\left(1+\frac{\lambda}{2}\right)\tilde{L}_{+}^{L}-\frac{\lambda}{2}\tilde{L}_{-}^{L}+a_{n},$
$\tilde{L}_{-}^{R}$ $=\frac{\lambda}{2}\tilde{L}_{+}^{L}+\left(1-\frac{\lambda}{2}\right)\tilde{L}_{-}^{L}+a_{n}.$ (34)
Thus the interface acts as a universal $\mathcal{H}_{L}\rightarrow\mathcal{H}_{R}$ map of the form $\mathcal{C}\circ\widetilde{\mathcal{S}}$ mapping the energy fluxes on left side to those on the right side with
$\widetilde{\mathcal{S}}=\begin{pmatrix}1+\frac{\lambda}{2}&-\frac{\lambda}{2}\\ \frac{\lambda}{2}&1-\frac{\lambda}{2}\end{pmatrix}$ (35)
being an invertible matrix and $\mathcal{C}$ is a conformal transformation, determined by $a_{n}$, on the right side that acts in the same way on the left and right movers . We note that $\widetilde{\mathcal{S}}$ and and $\mathcal{C}$ separately preserve the conformal boundary condition at the interface.
For any value of $\lambda$ between 0 and $2$, the interface is topological when
$a_{n}=\frac{\lambda}{2}\left(\tilde{L}_{-}^{L}-\tilde{L}_{+}^{R}\right).$ (36)
In this case, as evident from (Decoding the string in terms of holographic quantum maps), $\tilde{L}_{\pm}^{R}=\tilde{L}_{\pm}^{L}$ so that we obtain perfect energy transmission. Similarly, the interface is always factorizing when
$a_{n}=\tilde{L}_{+}^{R}-\tilde{L}_{-}^{L}$ (37)
as in this case $\tilde{L}_{+}^{R,L}=\tilde{L}_{-}^{R,L}$ and we obtain perfect energy reflection. Thus remarkably, the stringy interface is completely tunable. As discussed in the End Matter, the averaged null energy condition (ANEC) is not violated as the additional conformal transformations in the quantum maps do not affect the total energy fluxes.
Conclusions:- In this letter, we have demonstrated that the gravitational two-way junction including its stringy degrees of freedom can be decoded in terms of quantum maps. In absence of non-trivial stringy modes, these maps reduce to frequency independent universal maps [20, 31, 32, 21] that are realized by a defect operator, while the stringy modes imply that these maps should be generalized by composing them with appropriate conformal transformations. The full quantum maps preserve the conformal boundary condition on the interface. Particularly, the energy scattering involves independent and equal transmissions/reflections from both sides composed with an energy redistribution among the incoming/outgoing fluxes realized by conformal transformations. We have demonstrated the universality of these quantum maps by showing that they work for arbitrary input fluxes and an arbitrary background state (instead of the vacuum) on both sides of the interface within the universal sector.
Our calculations up to the linear order in the fluxes reveal that the holographic interface operator is a modification of (1) taking the form
$(L_{n,+}^{L}-L_{-n,-}^{L})I_{L,R}=I_{L,R}(\tilde{L}_{n,+}^{R}-\tilde{L}_{-n,-}^{R}),$ (38)
where $\tilde{L}_{\pm,m}$ are automorphisms of $L_{\pm,m}$ which preserve the Virasoro algebra. The understanding of the full non-linear gravitational problem, involving mixing of modes with different frequencies would be necessary to check if interface operators of the form (38) describe dual holographic junctions with stringy excitations.
It would be also of profound interest to understand how explicit reconstruction of sub-regions of the gravitational junction from interface CFT sub-regions can work by generalizing the reformulation of bulk reconstruction in holography in terms of recovery maps of quantum error correcting codes [4, 7]. As a first step, it would be of interest to understand how stringy degrees of freedom of gravitational junctions influence entanglement in the dual CFT following [33, 34, 35].
It would also be interesting to investigate the quantum null energy condition (QNEC) [36, 37, 38] in the holographic interfaces dual to gravitational junctions. The QNEC can impose non-trivial quantum thermodynamic bounds even if matter localized on bulk hypersurfaces satisfies the classical null energy condition [39, 40, 41] (see also [42] for a similar study in CFTs). The understanding of the link between quantum thermodynamic bounds and energy conditions is fundamental for the realization of quantum devices using interfaces in many-body systems.
Acknowledgements.
Acknowledgments:- We thank Costas Bachas, Avik Banerjee and Giuseppe Policastro for valuable discussions. We especially thank Giuseppe Policastro for discussions related to the ANEC bounds and quantum maps. AC, AM and MM acknowledge support from FONDECYT postdoctoral grant no. 3230222, FONDECYT regular grant no. 1240955 and “Doctorado Nacional” grant no. 21250596 of La Agencia Nacional de Investigación y Desarrollo (ANID), Chile, respectively. AM gratefully acknowledges the hospitality of LPENS, where a substantial part of this work was carried out during his tenure as a CNRS invited professor.
End matter
.1 Solution to the junction conditions
The solution to the junction conditions at linear order is below. Note that in the following solution we have turned off additional terms that solve the homogeneous equations, since they are not of the plane-wave form and don’t affect the scattering.
$\sigma_{d}=\epsilon\sigma e^{i\omega\tau}\left(\frac{(L^{L}_{-}-L^{R}_{+})e^{-i\omega\sigma p_{\lambda}}+(L^{L}_{+}-L^{R}_{-})e^{i\omega\sigma p_{\lambda}}}{2\omega^{2}}+\frac{i\lambda\left(4-\lambda^{2}\right)^{3/4}\left(A_{1}\cos\left(\frac{2\sigma\omega}{\sqrt{4-\lambda^{2}}}\right)-A_{2}\sin\left(\frac{2\sigma\omega}{\sqrt{4-\lambda^{2}}}\right)\right)}{4\sqrt{\pi}\sqrt{\omega}}\right),$ (39)
$\tau_{d}=\epsilon e^{i\omega\tau}\left(i\frac{(L^{L}_{-}-L^{R}_{+})e^{-i\omega\sigma p_{\lambda}}+(L^{L}_{+}-L^{R}_{-})e^{i\omega\sigma p_{\lambda}}}{4\omega^{3}(-2+\sigma^{2}\omega^{2})^{-1}}+\frac{\lambda\left(4-\lambda^{2}\right)^{3/4}\left(A_{1}\cos\left(\frac{2\sigma\omega}{\sqrt{4-\lambda^{2}}}\right)-A_{2}\sin\left(\frac{2\sigma\omega}{\sqrt{4-\lambda^{2}}}\right)\right)}{4\sqrt{\pi}\omega^{3/2}}\right),$ (40)
and $x_{d}=-\frac{\lambda\sigma}{\sqrt{4-\lambda^{2}}}+\epsilon x_{d}^{(1)}$, with
$x_{d}^{(1)}=e^{i\omega\tau}\left(\frac{-2\lambda\sigma\left((L^{L}_{-}+L^{R}_{+})e^{-i\omega\sigma p_{\lambda}}+(L^{L}_{+}+L^{R}_{-})e^{i\omega\sigma p_{\lambda}}\right)}{\sqrt{4-\lambda^{2}}\omega^{2}}+\frac{i\left((L^{L}_{-}+L^{R}_{+})e^{-i\omega\sigma p_{\lambda}}-(L^{L}_{+}+L^{R}_{-})e^{i\omega\sigma p_{\lambda}}\right)}{\omega^{3}(2+\sigma^{2}\omega^{2})^{-1}}\right),$ (41)
where $p_{\lambda}=\frac{\lambda}{\sqrt{4-\lambda^{2}}}$.
Ward identities at the interface and the displacement operator
To see the Ward identities we adopt continuous coordinates $\tilde{t},\tilde{x}$ where the energy-momentum tensors are
$\braket{\widetilde{T}^{L,R}_{\pm\pm}(\tilde{x}^{\pm})}=\frac{c\epsilon}{12\pi}e^{i\omega\tilde{x}_{2}^{\pm}}\tilde{L}_{\pm}^{L,R}+\mathcal{O}(\epsilon^{2})$ (42)
The full energy-momentum tensor can be written as
$\braket{\widetilde{T}_{\pm\pm}(\tilde{t},\tilde{x})}=\theta(-\tilde{x})\braket{\widetilde{T}^{L}_{\pm\pm}(\tilde{x}^{\pm})}+\theta(\tilde{x})\braket{\widetilde{T}^{R}_{\pm\pm}(\tilde{x}^{\pm})}$, (43)
where $\theta(x)$ is the Heaviside theta function. We then have the following Ward identities
$\partial_{\tilde{t}}\braket{\widetilde{T}^{\tilde{t}\tilde{t}}}+\partial_{\tilde{x}}\braket{\widetilde{T}^{\tilde{x}\tilde{t}}}=0,$ (44)
[…]
with the source
[…]
$=\frac{c\epsilon e^{i\omega\tilde{t}}}{12\pi}\left(\tilde{L}_{+}^{R}+\tilde{L}_{+}^{L}-\tilde{L}_{-}^{R}-\tilde{L}_{-}^{L}\right),$
$=\frac{c\epsilon e^{i\omega\tilde{t}}}{12\pi}\frac{4}{2+\lambda}(\tilde{L}_{+}^{R}-\tilde{L}_{-}^{L}-a_{n}).$ (46)
The Nambu-Goto mode appears in the source for the second Ward identity (45). The source for the first identity (44) vanishes because of energy conservation at the interface. The source $q(t)$ for the second Ward identity is the expectation value of the displacement operator $D$ (i.e. $q(t)=\langle D(t)\rangle)$ where $D$ is defined via
$D=2\left(\widetilde{T}_{++}^{R}-\widetilde{T}_{++}^{L}\right)=\left(\widetilde{T}_{++}^{R}-\widetilde{T}_{++}^{L}+\widetilde{T}_{--}^{R}-\widetilde{T}_{--}^{L}\right).$ (47)
The second equality above follows from energy conservation. The displacement operator quantifies the energy cost of a small displacement of the interface [43]. Eq. (46) indicates that the expectation value of the displacement operator is modified due to the presence of the Nambu-Goto mode.
Checking achronal ANEC
Finally, the achronal ANEC is always satisfied for the generalized quantum maps at the linear order as the linearized conformal transformations do not change the energy fluxes if $a_{n}(\omega)$ vanishes as $\omega\rightarrow 0$ (see (32) as for instance). Under the following transformation
$g(x^{\pm})=x^{\pm}+\epsilon\mathbb{g}(x^{\pm}),$ (48)
the averaged null energy
$\int{\rm d}x^{\pm}T_{\pm\pm}$ (49)
is invariant at $\mathcal{O}(\epsilon)$ if
$\int{\rm d}x^{\pm}\mathbb{g}^{\prime\prime\prime}(x^{+})=0,$ (50)
when $T_{\pm\pm}$ are also $\mathcal{O}(\epsilon)$. For $\mathbb{t}_{d}$ and other variables to be well defined, we actually require that $a_{n}(\omega)$ vanishes at least as $\omega^{3}$ as $\omega\rightarrow 0$.